# Description length of canonical and microcanonical models

### v3

@@ IV.2 Enforcing $n$ local constraints
$P_{\text{mic}}(\mathbf{G};\mathbf{r})=\begin{cases}\prod_{i=1}^{n}\frac{1}{\binom{m}{r_{i}}}&\text{if}\>\mathbf{r}(\mathbf{G})=\mathbf{r}\\ 0&\text{else}\end{cases}$ (56)
and the related maximum log-likelihood is the sum of the individual ones, i.e.
@@
$\text{COMP}_{\textrm{mic}}=n\ln[m+1]$ (58)
@@
$P_{\text{can}}(\mathbf{G};\bm{\theta})=\prod_{i=1}^{n}\frac{e^{-\theta_{i}r_{i}(\mathbf{G})}}{(1+e^{-\theta_{i}})^{m}}$ (60)
and can be re-written as
@@
$\text{DL}_{\text{can}}(\mathbf{G}^{*})=m\sum_{i=1}^{n}h\left(\frac{r_{i}}{m}\right)+n\ln\left[\frac{e^{m}\Gamma(m,m)}{m^{m-1}}+1\right].$ (65)
In the hypothesis of $m$ growing linearly with $n$, i.e.,$m=\Theta(n)$, we consider the following asymptotic expressions (for a derivation, see section S1.2 of the supplementary material)
@@
$a$ $=\frac{2}{3}\sqrt{\frac{2}{\pi}}\simeq 0.53,$ (68)
$b$ $=-\frac{11}{12}-\frac{4}{9\pi}\simeq-1.06.$ (69)
@@
As for the case of one global constraint, we focus on two aspects of these results. First, the difference between log-likelihood functions is either $\Theta(n)$ or $\Theta(n\ln n)$ depending on the order of $m$, i.e.,$m=\Theta(1)$ in the first case and $m=\Theta(n)$ in the second case. In any case, according to (36), the ensemble equivalence is broken. In particular, if $m=\Theta(n)$, the leading terms of expressions (IV.2) and (IV.2) are both of order $\Theta(n\ln n)$, thus cancelling out when compared - precisely as in the equivalent case. Nonetheless, the remaining terms are still of order $\Theta(n)$. Hence, the (asymptotic) difference between description lengths increases linearly with the system size, regardless of the growth rate of $m$.
@@ IV.3 Future research: Towards model-level equivalence
At this point, it is important to emphasize that comparing canonical and microcanonical ensembles is not the same as comparing the correspondent models. The former involves comparing two probability distributions, each obtained for specific parameter values, whereas the latter involves comparing two sets of distributions (i.e., two models), with each set sharing a common functional form within itself. While, in the context of the measure-level definition of ensemble equivalence, the comparison between ensembles can be reduced to comparing their likelihood terms, the comparison between models must take into account both their likelihoods and complexities.
@@
$\Delta\text{DL}$ might seem like a good candidate but is not suitable for this purpose: while it does take into account the model complexities, it still depends on the observed values. To be more general, we propose a definition involving the universal NML distributions representing the two models. Inspired by the measure level definition of non equivalence [6] of equation (35), we consider the following KL divergence
$D^{\text{NML}}_{\text{KL}}$ $\equiv S(P_{\text{mic}}^{\text{NML}}||P_{\text{can}}^{\text{NML}})$ (71)
@@
$=\overline{\Delta\text{DL}(\mathbf{c})},$
@@ V NML-based VS Bayesian approach to model selection
[…]
@@
$\text{DL}^{\textrm{NML}}_{\mathcal{M}}(\bm{x})=$ $-\ln\mathcal{L}_{\mathcal{M}}(\bm{x})+\frac{k}{2}\ln\frac{V}{2\pi}$
[…]
$\text{DL}_{\mathcal{M}}^{\text{Bayes}}(\bm{x})=$ $-\ln\mathcal{L}_{\mathcal{M}}(\bm{x})+\frac{k}{2}\ln\frac{V}{2\pi}$
$-\ln w(\hat{\bm{\theta}}(\bm{x}))+\ln\sqrt{\text{det}\mathbf{I}(\hat{\bm{\theta}}(\bm{x}))}+o(1),$ (79)
@@
[…]
by posing $w(\bm{\theta})\equiv J(\bm{\theta})$, the two asymptotic expressions coincide (for a formal derivation of these results, we redirect the interested reader to  [21] and  [18]). Thus, whenever the asymptotic formulas (V) and (V) hold true, the NML-based description length is asymptotically equivalent to the one provided by the Bayesian distribution equipped with Jeffreys prior (hereby, Bayes-Jeffreys description), which is NML-optimal.
@@
(see section S2.1 of the supplementary material for a derivation). In the limit $n\to\infty$ one finds
$\text{DL}_{\text{can}}^{\text{NML}}(\mathbf{G}^{*})-\text{DL}_{\text{can}}^{\text{Bayes-Jeffreys}}(\mathbf{G}^{*})=\frac{2}{3}\sqrt{\frac{2}{\pi nm}}+O\left(\frac{1}{nm}\right).$ (82)
@@
$\text{DL}_{\text{can}}^{\text{NML}}(\mathbf{G}^{*})-\text{DL}_{\text{can}}^{\text{Bayes-Jeffreys}}(\mathbf{G}^{*})=\frac{2}{3}\frac{\sqrt{2}n}{\sqrt{\pi m}}+\Theta\left(\frac{n}{m}\right).$ (84)
@@
stating that the NML-based description length of a canonical model coincides with the Bayesian description length of its microcanonical variant, equipped with the so-called canonical prior, introduced by Rissanen in [30] (see the definition in section S2.2 of the supplementary material). Moreover,
@@
The two identities, derived in S2.2, show that there exists a prior guaranteeing that the Bayesian description length of the canonical (microcanonical) model coincides with the NML-based description length of the microcanonical (canonical) model. This might lead to the conclusion that, in the framework of model selection, canonical and microcanonical models coincide even when non-equivalence holds if the right priors are chosen. Nevertheless, these priors are not NML-optimal. While this non-optimality is negligible in the case of one global constraint, it could play a crucial role when local constraints are enforced. In other words, in this case, different choices of prior can result in very different Bayesian description lengths for microcanonical as for canonical models.
The $n$ local constraints case examined in the previous sections provides an instructive example. In this case, we showed that the difference between the canonical and microcanonical NML description lengths is $\Theta(n)$. Combining this result with (85), we can conclude that when it comes to the microcanonical model, choosing between the uniform prior (which is NML-optimal) and the canonical prior can amount to a difference in description length of $\Theta(n)$ nats. Similarly, when it comes to the canonical model, it can be easily shown that, based on (84, 86) when $m=\Theta(n)$, choosing between the uniform and the Jeffreys prior can lead to a difference in description length of $\Theta(\sqrt{n})$ nats. In contrast, the same differences are $\Theta(1)$ when one global constraint is involved.
@@ VI Conclusion
We found that, although microcanonical models fit the data better than their canonical counterparts, they are also more complex. Determining the best-scoring model - i.e., the one that minimizes the description length - is not straightforward, as it depends on the observed values of the sufficient statistics. Specifically, we found that the canonical model performs best when its fit to the observed data exceeds its average fit across all possible values of the sufficient statistics.
@@ Acknowledgements
This work is supported by the European Union - NextGenerationEU - National Recovery and Resilience Plan (Piano Nazionale di Ripresa e Resilienza, PNRR), project ‘SoBigData.it - Strengthening the Italian RI for Social Mining and Big Data Analytics’ - Grant IR0000013 (n. 3264, 28/12/2021) (https://pnrr.sobigdata.it/).
Supplementary Materials
S1 Asymptotic Expansions
@@
$\ln\mathcal{L}_{\text{mic}}(l)=-\ln\binom{nm}{l}=-\ln(nm)!+\ln(nm-l)!+\ln l!.$ (S1)
@@
$\ln\mathcal{L}_{\text{mic}}(l)=-nm\ln(nm)+(nm-l)\ln(nm-l)+l\ln l-\frac{1}{2}\ln\left(\frac{nm}{2\pi l(nm-l)}\right)+o(1),$ (S3)
@@
$\ln\mathcal{L}_{\text{mic}}(p)=-nm\cdot h(p)+\frac{1}{2}\ln(nm)+\frac{1}{2}\ln(2\pi p(1-p))+o(1).$ (S4)
@@
$\ln\mathcal{L}_{\text{can}}(p)=-nm\cdot h(p),$ (S5)
@@
$\text{COMP}_{\text{can}}=\frac{1}{2}\ln(nm)+\frac{1}{2}\ln\left(\frac{\pi}{2}\right)+o(1).$ (S7)
@@
$\text{COMP}_{\textrm{can}}^{\text{NML}}\simeq\ln\left[\sqrt{\frac{\pi nm}{2}}+\frac{2}{3}+\frac{\sqrt{2\pi}}{24\sqrt{nm}}-\frac{4}{135nm}+\frac{\sqrt{2\pi}}{576\sqrt{(nm)^{3}}}+\frac{8}{2835(nm)^{2}}\right],$ (S8)
provided in  [25], where the authors show it to be very precise already for small values of the total number of entries $nm$. Finally, equation (54) is obtained by comparing the two asymptotic complexities. Notice that the asymptotic formula (S7) for the canonical complexity represents a well-known result, as it can be derived directly by applying the asymptotic formula (V) to the case of i.i.d. Bernoulli random variables.
S1.2 One-sided local constraints
@@
$\ln\mathcal{L}_{\textrm{mic}}(\mathbf{r})=-\sum_{i=1}^{n}\ln\binom{m}{r_{i}}=-n\ln m!+\sum_{i=1}^{n}\ln(m-r_{i})!+\ln r_{i}!.$ (S9)
@@
$\ln x!=x\ln x-x+\frac{1}{2}\ln(2\pi x)+\frac{1}{12x}+\Theta(1/x^{3}).$ (S10)
@@
$\ln\mathcal{L}_{\textrm{mic}}(\mathbf{r})=-m\sum_{i=1}^{n}h(p_{i})+\frac{n}{2}\ln m+\frac{1}{2}\sum_{i=1}^{n}[\ln(2\pi p_{i}(1-p_{i}))]\\ -\frac{n}{12m}+\frac{1}{12m}\sum_{i=1}^{n}\left[\frac{1}{p_{i}(1-p_{i})}\right]+o(1).$ (S11)
@@
$\ln\mathcal{L}_{\text{can}}(\mathbf{p})=-m\sum_{i=1}^{n}h(p_{i}),$ (S12)
@@
$=n\ln m+\frac{n}{m}+o(1),$ (S13)
which follows from $\ln\left(1+\frac{1}{x}\right)=\frac{1}{x}+\Theta(1/x^{2})$, is compared to the following asymptotic expansion of the canonical complexity, based on (S8)
$\text{COMP}_{\text{can}}=\frac{n}{2}\ln m+\frac{n}{2}\ln\frac{\pi}{2}+\frac{2}{3}\sqrt{\frac{2}{\pi}}\cdot\frac{n}{\sqrt{m}}+\left(\frac{1}{12}-\frac{4}{9\pi}\right)\cdot\frac{n}{m}+o(1)$ (S14)
@@
S2 NML and Bayes
S2.1 Bayesian-Jeffreys description lengths
In what follows, we compute the Bayesian description length $\text{DL}^{\text{B, Jeffreys}}_{\text{can}}$ of the canonical model obtained by constraining the sum $l$ over the matrix. As already stated, this model is equivalent to modelling $nm$ i.i.d. Bernoulli variable, and this result can be found in Example 8.3 of [18]. In the paper, we extend this result to the case of $n$ one-sided local constraints.

### v4

@@ IV.2 Enforcing $n$ local constraints
$P_{\text{mic}}(\mathbf{G};\mathbf{r})=\begin{cases}\prod_{i=1}^{n}\frac{1}{\binom{m}{r_{i}}}&\text{if}\>\mathbf{r}(\mathbf{G})=\mathbf{r}\\ 0&\text{else,}\end{cases}$ (56)
and the related maximum log-likelihood is the sum of the individual ones, i.e.,
@@
$\text{COMP}_{\textrm{mic}}=n\ln[m+1],$ (58)
@@
$P_{\text{can}}(\mathbf{G};\bm{\theta})=\prod_{i=1}^{n}\frac{e^{-\theta_{i}r_{i}(\mathbf{G})}}{(1+e^{-\theta_{i}})^{m}},$ (60)
and can be rewritten as
@@
$\text{DL}_{\text{can}}(\mathbf{G}^{*})=m\sum_{i=1}^{n}h\left(\frac{r_{i}}{m}\right)\\ +n\ln\left[\frac{e^{m}\Gamma(m,m)}{m^{m-1}}+1\right].$ (65)
In the hypothesis of $m$ growing linearly with $n$, i.e., $m=\Theta(n)$, we consider the following asymptotic expressions (for a derivation, see section B.2 of the Appendix)
@@
$a$ $=\frac{2}{3}\sqrt{\frac{2}{\pi}}\simeq 0.53,$
$b$ $=-\frac{11}{12}-\frac{4}{9\pi}\simeq-1.06.$
@@
As for the case of one global constraint, we focus on two aspects of these results. First, the difference between log-likelihood functions is either $\Theta(n)$ or $\Theta(n\ln n)$ depending on the order of $m$, i.e., $m=\Theta(1)$ in the first case and $m=\Theta(n)$ in the second case. In any case, according to (36), the ensemble equivalence is broken. In particular, if $m=\Theta(n)$, the leading terms of expressions (IV.2) and (IV.2) are both of order $\Theta(n\ln n)$, thus canceling out when compared, precisely as in the equivalent case. Nonetheless, the remaining terms are still of order $\Theta(n)$. Hence, the (asymptotic) difference between description lengths increases linearly with the system size, regardless of the growth rate of $m$.
@@ IV.3 Future research: Towards model-level equivalence
At this point, it is important to emphasize that comparing canonical and microcanonical ensembles is not the same as comparing the corresponding models. The former involves comparing two probability distributions, each obtained for specific parameter values, whereas the latter involves comparing two sets of distributions (i.e., two models), with each set sharing a common functional form within itself. While in the context of the measure-level definition of ensemble equivalence, the comparison between ensembles can be reduced to comparing their likelihood terms, the comparison between models must take into account both their likelihoods and complexities.
@@
$\Delta\text{DL}$ might seem like a good candidate but is not suitable for this purpose: while it does take into account the model complexities, it still depends on the observed values. To be more general, we propose a definition involving the universal NML distributions representing the two models. Inspired by the measure level definition of non-equivalence [18] of equation (35), we consider the following KL divergence
$D^{\text{NML}}_{\text{KL}}$ $\equiv S(P_{\text{mic}}^{\text{NML}}||P_{\text{can}}^{\text{NML}})$
@@
$=\overline{\Delta\text{DL}(\mathbf{c})},$ (69)
@@ V NML-based VS Bayesian approach to model selection
$\bar{P}^{\text{Bayes}}_{\mathcal{M}}(\bm{x})=\int_{\Theta}P_{\mathcal{M}}(\bm{x};\bm{\theta})w(\bm{\theta})d\bm{\theta}$ (72)
@@
$\text{DL}^{\textrm{NML}}_{\mathcal{M}}(\bm{x})=-\ln\mathcal{L}_{\mathcal{M}}(\bm{x})+\frac{k}{2}\ln\frac{V}{2\pi}\\ +\ln\int_{\Theta}\sqrt{\text{det}\mathbf{I}(\bm{\theta})}d\bm{\theta}+o(1),$ (76)
$\text{DL}_{\mathcal{M}}^{\text{Bayes}}(\bm{x})=-\ln\mathcal{L}_{\mathcal{M}}(\bm{x})+\frac{k}{2}\ln\frac{V}{2\pi}\\ -\ln w(\hat{\bm{\theta}}(\bm{x}))+\ln\sqrt{\text{det}\mathbf{I}(\hat{\bm{\theta}}(\bm{x}))}+o(1),$ (77)
@@
$J(\bm{\theta})=\frac{\sqrt{\text{det}\mathbf{I}(\bm{\theta})}}{\int_{\Theta}\sqrt{\text{det}\mathbf{I}(\bm{\theta})}d\bm{\theta}}$ (78)
by posing $w(\bm{\theta})\equiv J(\bm{\theta})$, the two asymptotic expressions coincide (for a formal derivation of these results, we redirect the interested reader to  [33] and  [30]). Thus, whenever the asymptotic formulas (76) and (77) hold true, the NML-based description length is asymptotically equivalent to the one provided by the Bayesian distribution equipped with Jeffreys prior (hereby, Bayes-Jeffreys description), which is NML-optimal.
@@
(see section C.1 of the Appendix for a derivation). In the limit $n\to\infty$ one finds
$\text{DL}_{\text{can}}^{\text{NML}}(\mathbf{G}^{*})-\text{DL}_{\text{can}}^{\text{Bayes-Jeffreys}}(\mathbf{G}^{*})=\\ =\frac{2}{3}\sqrt{\frac{2}{\pi nm}}+\Theta\left(\frac{1}{nm}\right).$ (80)
@@
$\text{DL}_{\text{can}}^{\text{NML}}(\mathbf{G}^{*})-\text{DL}_{\text{can}}^{\text{Bayes-Jeffreys}}(\mathbf{G}^{*})=\\ =\frac{2}{3}\frac{\sqrt{2}n}{\sqrt{\pi m}}+\Theta\left(\frac{n}{m}\right).$ (82)
@@
stating that the NML-based description length of a canonical model coincides with the Bayesian description length of its microcanonical variant, equipped with the so-called canonical prior, introduced by Rissanen in [44] (see the definition in section C.2 of the Appendix). Moreover,
@@
The two identities, derived in C.2, show that there exists a prior guaranteeing that the Bayesian description length of the canonical (microcanonical) model coincides with the NML-based description length of the microcanonical (canonical) model. This might lead to the conclusion that, in the framework of model selection, canonical and microcanonical models coincide even when non-equivalence holds if the right priors are chosen. Nevertheless, these priors are not NML-optimal. While this non-optimality is negligible in the case of one global constraint, it could play a crucial role when local constraints are enforced. In other words, in this case, different choices of prior can result in very different Bayesian description lengths for microcanonical as for canonical models.
The $n$ local constraints case examined in the previous sections provides an instructive example. In this case, we showed that the difference between the canonical and microcanonical NML description lengths is $\Theta(n)$. Combining this result with (83), we can conclude that when it comes to the microcanonical model, choosing between the uniform prior (which is NML-optimal) and the canonical prior can amount to a difference in description length of $\Theta(n)$ nats. Similarly, when it comes to the canonical model, it can be easily shown that, based on (82, 84) when $m=\Theta(n)$, choosing between the uniform and the Jeffreys prior can lead to a difference in description length of $\Theta(\sqrt{n})$ nats. In contrast, the same differences are $\Theta(1)$ when one global constraint is involved.
@@ VI Conclusion
We found that, although microcanonical models fit the data better than their canonical counterparts, they are always more complex. Determining the best-scoring model - i.e., the one that minimizes the description length - is not straightforward, as it depends on the observed values of the sufficient statistics. Specifically, we found that the canonical model performs best when its fit to the observed data exceeds its average fit across all possible values of the sufficient statistics.
@@ Acknowledgements
This work is supported by the European Union - NextGenerationEU - National Recovery and Resilience Plan (Piano Nazionale di Ripresa e Resilienza, PNRR), projects ‘SoBigData.it - Strengthening the Italian RI for Social Mining and Big Data Analytics’ - Grant IR0000013 (n. 3264, 28/12/2021) (https://pnrr.sobigdata.it/); and “Reconstruction, Resilience and Recovery of Socio-Economic Networks” RECON-NET EP_FAIR_005 - PE0000013 “FAIR” - PNRR M4C2 Investment 1.3.
Appendix A Derivation of the microcanonical distribution from entropy maximization
It is well known that canonical maximum-entropy models can be derived by maximizing the Shannon entropy subject to soft constraints [35]. Here, we show that microcanonical models can also be seen as the solution to an entropy maximization problem under hard constraints, as a direct consequence of the Shannon-Khinchin axioms that uniquely identify Shannon entropy [45, 46].
Let $\mathcal{G}$ denote the set of all configurations, and let $\mathcal{G}_{\text{valid}}\subset\mathcal{G}$ denote the subset of configurations that exactly satisfy the hard constraints. Consider the problem of maximizing the Shannon entropy
$S[P]=-\sum_{\mathbf{G}\in\mathcal{G}}P(\mathbf{G})\log P(\mathbf{G})$ (85)
over probability distributions $P(\mathbf{G})$ supported on $\mathcal{G}$, under the requirement that $P(\mathbf{G})=0$ for all $\mathbf{G}\notin\mathcal{G}_{\text{valid}}$. That is, configurations violating the constraints are assigned zero probability.
The third Shannon-Khinchin axiom states that $S[P]$ is expansible, i.e., it does not change if an outcome with zero probability is added. This implies that, in this case, maximizing entropy over the full space $\mathcal{G}$ with support restricted to $\mathcal{G}_{\text{valid}}$ is equivalent to maximizing entropy over $\mathcal{G}_{\text{valid}}$ only. Then, according to the second Shannon-Khinchin axiom, which states that entropy is maximized by the uniform distribution in the absence of further constraints, the solution is the uniform distribution over $\mathcal{G}_{\text{valid}}$:
$P_{\text{mic}}(\bm{x})=\begin{cases}\frac{1}{\left\lvert\mathcal{G}_{\text{valid}}\right\rvert}&\text{if }\mathbf{G}\in\mathcal{G}_{\text{valid}},\\ 0&\text{else}.\end{cases}$ (86)
With this derivation, canonical and microcanonical ensembles emerge as distinct solutions to the same entropy maximization principle, differing only in how the constraints are imposed.
Appendix B Asymptotic expansions
@@ B.1 One global constraint
$\ln\mathcal{L}_{\text{mic}}(l)$ $=-\ln\binom{nm}{l}$
$=-\ln(nm)!+\ln(nm-l)!+\ln l!.$ (87)
@@
$\ln\mathcal{L}_{\text{mic}}(l)=-nm\ln(nm)+(nm-l)\ln(nm-l)\\ +l\ln l-\frac{1}{2}\ln\left(\frac{nm}{2\pi l(nm-l)}\right)+o(1),$ (89)
@@
$\ln\mathcal{L}_{\text{mic}}(p)=-nm\cdot h(p)+\frac{1}{2}\ln(nm)\\ +\frac{1}{2}\ln(2\pi p(1-p))+o(1).$ (90)
@@
$\ln\mathcal{L}_{\text{can}}(p)=-nm\cdot h(p),$ (91)
@@
$\text{COMP}_{\text{can}}=\frac{1}{2}\ln(nm)+\frac{1}{2}\ln\left(\frac{\pi}{2}\right)+o(1).$ (93)
@@
$\text{COMP}_{\textrm{can}}^{\text{NML}}\simeq\ln\left[\sqrt{\frac{\pi nm}{2}}+\frac{2}{3}+\frac{\sqrt{2\pi}}{24\sqrt{nm}}-\frac{4}{135nm}\right.\\ +\left.\frac{\sqrt{2\pi}}{576\sqrt{(nm)^{3}}}+\frac{8}{2835(nm)^{2}}\right],$ (94)
provided in [39], where the authors show it to be very precise already for small values of the total number of entries $nm$. Finally, equation (54) is obtained by comparing the two asymptotic complexities. Notice that the asymptotic formula (93) for the canonical complexity represents a well-known result, as it can be derived directly by applying the asymptotic formula (76) to the case of i.i.d. Bernoulli random variables.
B.2 One-sided local constraints
@@
$\ln\mathcal{L}_{\textrm{mic}}(\mathbf{r})$ $=-\sum_{i=1}^{n}\ln\binom{m}{r_{i}}$
$=-n\ln m!+\sum_{i=1}^{n}\ln(m-r_{i})!+\ln r_{i}!.$ (95)
@@
$\ln x!=x\ln x-x+\frac{1}{2}\ln(2\pi x)+\frac{1}{12x}+\Theta(1/x^{3}).$ (96)
@@
$\ln\mathcal{L}_{\textrm{mic}}(\mathbf{r})=-m\sum_{i=1}^{n}h(p_{i})+\frac{n}{2}\ln m\\ +\frac{1}{2}\sum_{i=1}^{n}[\ln(2\pi p_{i}(1-p_{i}))]-\frac{n}{12m}\\ +\frac{1}{12m}\sum_{i=1}^{n}\left[\frac{1}{p_{i}(1-p_{i})}\right]+o(1).$ (97)
@@
$\ln\mathcal{L}_{\text{can}}(\mathbf{p})=-m\sum_{i=1}^{n}h(p_{i}),$ (98)
@@
$=n\ln m+\frac{n}{m}+o(1),$ (99)
which follows from $\ln\left(1+\frac{1}{x}\right)=\frac{1}{x}+\Theta(1/x^{2})$, is compared to the following asymptotic expansion of the canonical complexity, based on (94)
$\text{COMP}_{\text{can}}=\frac{n}{2}\ln m+\frac{n}{2}\ln\frac{\pi}{2}+\frac{2}{3}\sqrt{\frac{2}{\pi}}\cdot\frac{n}{\sqrt{m}}\\ +\left(\frac{1}{12}-\frac{4}{9\pi}\right)\cdot\frac{n}{m}+o(1)$ (100)
@@
Appendix C NML and Bayes
C.1 Bayesian-Jeffreys description lengths
In what follows, we compute the Bayesian description length $\text{DL}^{\text{B, Jeffreys}}_{\text{can}}$ of the canonical model obtained by constraining the sum $l$ over the matrix. As already stated, this model is equivalent to modeling $nm$ i.i.d. Bernoulli variables, and this result can be found in Example 8.3 of [30]. In the paper, we extend this result to the case of $n$ one-sided local constraints.